OpenStudy (shyniah365):

Match each set of ordered pairs with their rise over run. Match Term Definition two comma negative three and negative one comma four A) negative seven‾thirds negative two comma three and negative one comma four B) one‾third two comma negative three and one comma four C) 1 negative two comma three and one comma four D) negative seven

OpenStudy (shyniah365):

i give medals

OpenStudy (texaschic101):

slope formula : slope = (y2 - y1) / (x2 - x1) (2,-3) x1 = 2 and y1 = -3 (-1, 4 ) x2 = -1 and y2 = 4 now we sub slope = (4 - (-3) / (-1 - 2) slope = (4 + 3) / -3 slope = -7/3 ===> A (-2,3) x1 = -2 and y1 = 3 (-1,4) x2 = -1 and y2 = 4 slope = (y2 - y1) / (x2 - x1) slope = (4 - 3) / (-1 - (-2) slope = 1/(-1 + 2) slope = 1/1 slope = 1 ====> B (2,-3) x1 = 2 and y1 = -3 (1,4) x2 = 1 and y2 = 4 slope = (y2 - y1) / (x2 - x1) slope = (4 - (-3) / (1 - 2) slope = (4 + 3) / -1 slope = -7 ====> D (-2,3) x1 = -2 and y1 = 3 (-1,4) x2 = -1 and y2 = 4 slope = (y2 - y1) / (x2 - x1) slope = (4 - 3) / (-1 - (-2) slope = 1/(-1 + 2) slope = 1 ====> C

undeadknight26 (undeadknight26):

@shyniah365 did u get your answers?

OpenStudy (shyniah365):

yes i did thank you for helping me